Grewett Publishers Inc. budgeted production of 54,000 diaries in 2016. Each diary requires assembly. Assume that five minutes are required to assemble each diary.
If assembly labor costs $13.75 per hour, determine the direct labor cost budget for 2016. Do not round your intermediate calculations but, if required, round your final answer to the nearest dollar.
Answer #1

Correct Answer :

The direct labour cost budget for 2016 is $ 61875.

Working:

A

budgeted production

54000

units

B

assembly time required by each unit

5

minutes

C

assembly cost per hour

$           13.75

dollar

D = A*B

total assembly time in minutes

270000

minutes

E = D/60

total assembly time in hours

4500

hours

F = E*C

total direct labour cost budget

$   61,875.00

dollars

Kindly give a comment if problem in understanding any part of the question.

consider giving positive review if possible.

Thanks in advance.